AntonyTheGreatest:

Find f(–2) for the function f(x) = 3x2 – 2x + 7. −13 −1 1 23

Isolation:

substitute -2 as x in \(3x^2-2x+7\) so do \[3(-2^2)-2(-2)+7\]

AntonyTheGreatest:

im trying to get te answer

Isolation:

use PEMDAS and simplify

Isolation:

first, what is -2 squared?

AntonyTheGreatest:

-4

Isolation:

no, -2 squared is 4

Isolation:

\(3(-2^2)-2(-2)+7\\3(4)-2(-2)+7\)

Isolation:

and simplify from there

Isolation:

what is 3 times 4, what is -2 times -2

AntonyTheGreatest:

12

AntonyTheGreatest:

-2 x -2 =4

Isolation:

@isolation wrote:
\(3(-2^2)-2(-2)+7\\3(4)-2(-2)+7\)
\(12+4+7\)

Isolation:

then add

AntonyTheGreatest:

12+4+7 12+4=16+7=23

Isolation:

yes so there's your answer

AntonyTheGreatest:

you reaady for the next one

Isolation:

sure post it